# Exact Random Graph Matching with Multiple Graphs

### v1

@@ 4.2 Achievability of exact graph matching (Theorem 3)
    Each instance of pairwise matching using the $k$-core estimator is correct on its domain, i.e.
    $\widehat{\mu}_{ij}(v)=\pi^{*}_{ij}(v)\ \forall v\in\mathsf{dom}(\widehat{\mu}_{ij}),\ \forall i,j.$
@@
On step 1
This falls back to the regime of analyzing the performance of the $k$-core estimator in the setting of two graphs. Cullina and co-authors [CKMP19] showed that the $k$-core estimator is precise: For any two correlated graphs $G_{i}$ and $G_{j}$ with $p=C\log(n)/n$ and constant $s$, the $k$-core estimator correctly matches all nodes in $\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$ with probability $1-o(1)$. In fact, this is true for any $C>0$ and for any $k\geq 13$ [RS23]. Therefore, using the fact that the number of instances of pairwise matchings is constant whenever $m$ is constant, a union bound reveals
@@
$\leq\sum_{i=1}^{m}\sum_{j=1}^{m}\mathbb{P}\left(\widehat{\mu}_{i,j}(v)\neq\pi^{*}_{i,j}(v)\text{ for some }v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right)$
$=o(1).$
We have proved the following.
Proposition 5.
Let $G_{1},\cdots,G_{m}$ be correlated graphs from the subsampling model. Let $k\geq 13$ and let $\widehat{\mu}_{ij}$ denote the matching output by the $k$-core estimator on graphs $G_{i}$ and $G_{j}$. Then,
( $\mathbb{P}(\exists\ 1\leq i<j\leq m,\text{ and }v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime}))\text{ such that }\widehat{\mu}_{ij}(v)\neq\pi^{*}_{ij}(v))=o(1).$
On step 2
@@
The challenging part of the proof is to show that boosting through transitive closure matches all the nodes with probability $1-o(1)$ if $Cs(1-(1-s)^{m-1})>1$. It is instructive to visualize this using transitivity graphs.
Definition 6 (Transitivity graph, $\mathcal{H}(v)$).
@@
For each node $v\in V$, let $\mathcal{H}(v)$ denote the graph on the vertex set $\{g_{1},\cdots,g_{m}\}$ such that an edge $\{g_{i},g_{j}\}$ is present in $\mathcal{H}(v)$ if and only if $v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$.
@@
On the event that each instance of pairwise matching using the $k$-core is correct, the edge $\{g_{i},g_{j}\}$ is present in $\mathcal{H}(v)$ if and only if $v$ is correctly matched using the $k$-core estimator between $G_{i}$ and $G_{j}$, i.e. $\pi^{*}_{1i}(v)$ is matched to $\pi^{*}_{1j}(v)$. Thus, in order for Step 2 to succeed (i.e. to exactly match all vertices across all graphs), it suffices that the graph $\mathcal{H}(v)$ is connected for each node $v\in V$. However, studying the connectivity of the transitivity graphs is challenging because in any graph $\mathcal{H}(v)$, no two edges are independent. This is because the $k$-cores of any two intersection graphs $G_{a}^{\prime}\wedge G_{b}^{\prime}$ and $G_{c}^{\prime}\wedge G_{d}^{\prime}$ are correlated, because all the graphs $G_{a},G_{b},G_{c}$ and $G_{d}$ are themselves correlated. To overcome this, we introduce another graph ~$\widetilde{\mathcal{H}}(v)$ that relates to $\mathcal{H}(v)$ and is amenable to analysis.
Definition 7.
@@
For each node $v\in V$, let ~$\widetilde{\mathcal{H}}(v)$ denote a complete weighted graph on the vertex set $\{g_{1},\cdots,g_{m}\}$ such that the weight on any edge $\{g_{i},g_{j}\}$ is $\widetilde{c}_{v}\left(i,j\right):=\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v).$
The relationship between the graphs $\mathcal{H}(v)$ and ~$\widetilde{\mathcal{H}}(v)$ stems from a useful relationship between the degree of node $v$ in $G_{i}^{\prime}\wedge G_{j}^{\prime}$ and the inclusion of $v$ in $\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$ for each $i$ and $j$. Since this result is of independent interest in the study of random graphs, we state it below for general Erdős-Rényi graphs.
Lemma 8.
Let $n$ and $k$ be positive integers and let $G\sim\mathsf{ER}(n,\alpha\log(n)/n)$ for some $\alpha>0$. Let $v$ be a node of $G$ and let $\delta_{G}(v)$ denote the degree of $v$ in $G$. Then,
($\mathbb{P}\left(\left\{v\notin\mathsf{core}_{k}(G)\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\right)=o\left(1/n\right).$ (1)
@@
For any $i$ and $j$, the graph $G_{i}^{\prime}\wedge G_{j}^{\prime}\sim\mathsf{ER}(n,Cs^{2}\log(n)/n)$. Thus, Lemma 8 implies that with probability $1-o(1/n)$, if a pair $\{g_{i},g_{j}\}$ has edge weight $\widetilde{c}_{ij}\geq k+1/\alpha$ in ~$\widetilde{\mathcal{H}}(v)$, then the corresponding edge $\{g_{i},g_{j}\}$ is present in the transitivity graph $\mathcal{H}(v)$. Equivalently, $v$ is correctly matched between $G_{i}$ and $G_{j}$ in the instance of pairwise $k$-core matching between them.
The graph $\mathcal{H}(v)$ is not connected only if it contains a (non-empty) vertex cut $U\subset\{1,\cdots,m\}$ with no edge crossing between $U$ and $U^{c}$. Let $c_{v}(U)$ denote the number of such crossing edges in $\mathcal{H}(v)$. Furthermore, define the cost of the cut $U$ in ~$\widetilde{\mathcal{H}}(v)$ as
@@
Lemma 8 is a statement about a single graph, but we show it can be invoked to prove the following.
Theorem 9.
Let $G_{1},\cdots,G_{m}$ be correlated graphs from the subsampling model with parameters $C$ and $s$. Let $v\in V$ and let $U$ be a vertex cut of $\{1,\cdots,m\}$ such that $|U|\leq\lfloor m/2\rfloor$. Then,
($\mathbb{P}\left(\left\{c_{v}(U)=0\right\}\cap\left\{\widetilde{c}_{v}(U)>\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)\right\}\right)=o(1/n).$ (2)
It suffices therefore to analyze the probability that the graph ~$\widetilde{\mathcal{H}}(v)$ has a cut $U$ such that its cost $\widetilde{c}_{v}(U)$ is too small. To that end, we show that the bottleneck arises from vertex cuts of small size. Formally,
Theorem 10.
Let $G_{1},\cdots,G_{m}$ be correlated graphs from the subsampling model. Let $v\in V$ and let $U_{\ell}$ denote the set $\{1,\cdots,\ell\}$ for $\ell$ in $\{1,\cdots,\lfloor m/2\rfloor\}$. For any vertex cut $U$ of $\{1,\cdots,m\}$, let $\widetilde{c}_{v}(U)$ denote its cost in the graph ~$\widetilde{\mathcal{H}}(v)$. The following stochastic ordering holds:
@@
Theorems 9 and 10 imply that the tightest bottleneck to the connectivity of $\mathcal{H}(v)$ is the event that $\widetilde{c}_{v}(U_{1})$ is below the threshold $r:=\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)$, i.e. the sum of degrees of $v$ over the intersection graphs $(G_{1}\wedge G_{j}^{\prime}:j=2,\cdots,m)$ is less than $r$. This event occurs only if the degree of $v$ is less than $r$ in each of the intersection graphs $(G_{1}\wedge G_{j}^{\prime}:j=2,\cdots,m)$. However, under the condition $Cs(1-(1-s)^{m-1})>1$, it turns out that this event occurs with probability $o(1/n)$.
Theorem 11.
@@
Let $G_{1},\cdots,G_{m}$ be obtained from the subsampling model with parameters $C$ and $s$. Let $r=\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)$. Let $v\in[n]$ and suppose that $Cs(1-(1-s)^{m-1})>1$. Then,
@@
$\mathbb{P}\left(\widetilde{c}_{v}(U_{1})\leq r\right)\leq\mathbb{P}\left(\left\{\delta_{G_{1}\wedge G_{2}^{\prime}}(v)\leq r\right\}\cap\cdots\cap\left\{\delta_{G_{1}\wedge G_{m}^{\prime}}(v)\leq r\right\}\right)=o\left(1/n\right).$
4.3 Piecing it all together: Proof of Theorem 3
@@
Proof of Theorem 3. Let $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ denote the output of Algorithm 2 with $k\geq 13$. Let $E_{1}$ (resp. $E_{2}$) denote the event that Algorithm 1 (resp. Algorithm 2) fails to match all $m$ graphs exactly, i.e.
$E_{1}=\left\{\widehat{\pi}_{12}^{\text{ML}}\neq\pi^{*}_{12}\right\}\cup\cdots\cup\left\{\widehat{\pi}_{1m}^{\text{ML}}\neq\pi^{*}_{1m}\right\},\quad\quad E_{2}=\left\{\widehat{\pi}_{12}\ \neq\pi^{*}_{12}\right\}\cup\cdots\cup\left\{\widehat{\pi}_{1m}\neq\pi^{*}_{1m}\right\}.$
@@
First, we show that the output of Algorithm 2 is correct with probability $1-o(1)$ whenever $Cs(1-(1-s)^{m-1})>1$. If the event $E_{2}$ occurs, then either step 1 failed, i.e. there is a $k$-core matching $\widehat{\mu}_{ij}$ that is incorrect, or step 2 failed, i.e. at least one of the graphs $\mathcal{H}(v)$ is not connected. Therefore,
@@
$\mathbb{P}\left(E_{2}\right)$ $\leq\mathbb{P}\left(\bigcup_{i,j}\ \bigcup_{v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})}\!\!\!\!\left\{\widehat{\mu}_{ij}\neq\pi^{*}_{ij}\right\}\!\right)\!+\mathbb{P}\left(\bigcup_{v\in V}\left\{\mathcal{H}(v)\text{ is not connected}\right\}\!\right)\!\leq o(1)+\!\sum_{v\in V}q_{v},$
@@
where the last step uses Proposition 5, and $q_{v}$ denotes the probability that the transitivity graph $\mathcal{H}(v)$ is not connected. For each $\ell$ in the set $\{1,\cdots,\lfloor m/2\rfloor\}$, let $U_{\ell}$ denote the set $\left\{1,\cdots,\ell\right\}$. Then,
@@
$q_{v}\!$ $=\mathbb{P}\left(\bigcup_{\ell=1}^{\lfloor m/2\rfloor}\left\{\exists\ U\subset\{1,\cdots,m\}:|U|=\ell\text{ and }c_{v}(U)=0\right\}\right)$
@@
$\leq\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\cdot\mathbb{P}\left(c_{v}(U_{\ell})=0\right)$
@@
$\leq\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\left[\mathbb{P}\left(\widetilde{c}_{v}(U_{\ell})\leq\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\!\right)\!+\!\mathbb{P}\left(\left\{c_{v}(U_{\ell})=0\right\}\cap\bigg{\{}\widetilde{c}_{v}(U_{\ell})>\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\bigg{\}}\right)\right]$
@@
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\left[\mathbb{P}\left(\widetilde{c}_{v}(U_{\ell})\leq\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\!\right)\!+\!o\left(\frac{1}{n}\right)\right]$
@@
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\left[\mathbb{P}\left(\widetilde{c}_{v}(U_{1})\leq\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\!\right)\!+\!o\left(\frac{1}{n}\right)\right]$

### v2

@@ 5.1 Analyzing Algorithm 1: Intuition behind the proof of Theorem 1
    Each instance of pairwise matching using the $k$-core estimator is precise, i.e. it is correct on its domain with probability $1-o(1)$, i.e. $\widehat{\mu}_{ij}(v)=\pi^{*}_{ij}(v)$ for all $v\in\mathrm{dom}(\widehat{\mu}_{ij})$ for all $i$ and $j$.
@@
5.1.1 On step 1
This falls back to the regime of analyzing the performance of the $k$-core estimator in the setting of two graphs. Cullina and co-authors (CKMP, 19) showed that the $k$-core estimator is precise: For any two correlated Erdős-Rényi graphs $G_{i}$ and $G_{j}$ with $\mathbf{P}_{ij}=C\log(n)/n$ for all $1\leq i<j\leq n$, and constant subsampling probability $s$, the $k$-core estimator with sufficiently large $k$ correctly matches all nodes in $\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$ with probability $1-o(1)$. Building on this analysis, Racz and Sridhar (RS, 23) showed that this is true more generally:
Lemma 20 (Lemma III.4 from (RS, 23))
Let $G_{1}$ and $G_{2}$ be correlated inhomogeneous graphs obtained by subsampling each edge independently with probability $s$ from a parent graph $G\sim\mathrm{RG}(n,\mathbf{P})$. Suppose that there exists some $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}\leq o(n^{\alpha/2-\varepsilon-1})$, and let $k>\frac{12}{1-\alpha+2\varepsilon}$. Then, with probability $1-o(1)$, it holds that the matching $\widehat{\mu}_{k}$ output by the $k$-core estimator satisfies
$\mathrm{dom}(\widehat{\mu}_{k})=\mathrm{core}_{k}(G_{1}^{\prime}\wedge G_{2}^{\prime})\text{ and }\widehat{\mu}_{k}(v)=\pi^{*}(v)\ \forall\ v\in\mathrm{dom}(\widehat{\mu}_{k}).$
This precision of the $k$-core estimator is critical to the success of Algorithm 1 since it ensures that there are no conflicts during the boosting step. Applying Lemma 20 to the setting of $m$ graphs yields that each instance of pairwise matching in step $1$ of Algorithm 1 is indeed precise.
Corollary 21
Let $\widehat{\mu}_{ij}$ denote the matching output by the $k$-core estimator on graphs $G_{i}$ and $G_{j}$. Under the conditions of Lemma 20,
( $\mathbb{P}(\exists\ 1\leq i<j\leq m,\text{ and }v\in\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\text{ such that }\widehat{\mu}_{ij}(v)\neq\pi^{*}_{ij}(v))=o(1).$
Proof. Since the number of instances of pairwise matchings is constant whenever $m$ is constant, a union bound yields
( $\mathbb{P}(\exists\ 1$ $\leq i<j\leq m,\text{ and }v\in\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\text{ such that }\widehat{\mu}_{ij}(v)\neq\pi^{*}_{ij}(v))$
$\leq\sum_{i=1}^{m}\sum_{j=1}^{m}\mathbb{P}\left(\widehat{\mu}_{ij}(v)\neq\pi^{*}_{ij}(v)\text{ for some }v\in\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right)=o(1).$
@@ 5.1.1 On step 1
5.1.2 On step 2
@@
The challenging part of the proof is to show that boosting through transitive closure matches all the nodes with probability $1-o(1)$ under the sufficient condition. It is instructive to visualize this using transitivity graphs.
Definition 22 (Transitivity graph, $\mathcal{H}(v)$)
@@
For each node $v\in[n]$, let $\mathcal{H}(v)$ denote the graph on the vertex set $\{g_{1},\cdots,g_{m}\}$ such that an edge $\{g_{i},g_{j}\}$ is present in $\mathcal{H}(v)$ if and only if $v\in\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$.
@@
On the event that each instance of pairwise matching using the $k$-core is correct, the edge $\{g_{i},g_{j}\}$ is present in $\mathcal{H}(v)$ if and only if $v$ is correctly matched using the $k$-core estimator between $G_{i}$ and $G_{j}$, i.e. $\pi^{*}_{1i}(v)$ is matched to $\pi^{*}_{1j}(v)$. Thus, in order for step 2 to succeed (i.e. to exactly match all vertices across all graphs), it suffices that the graph $\mathcal{H}(v)$ be connected for each node $v\in[n]$. However, studying the connectivity of the transitivity graphs is challenging because in any graph $\mathcal{H}(v)$, no two edges are independent. This is because the $k$-cores of any two intersection graphs $G_{a}^{\prime}\wedge G_{b}^{\prime}$ and $G_{c}^{\prime}\wedge G_{d}^{\prime}$ are correlated, because all the graphs $G_{a},G_{b},G_{c}$ and $G_{d}$ are themselves correlated. To overcome this, we introduce another graph ~$\widetilde{\mathcal{H}}(v)$ that relates to $\mathcal{H}(v)$ and is amenable to analysis. Let $\delta_{G}(v)$ denote the degree of a node $v$ in graph $G$.
Definition 23
@@
For each node $v\in[n]$, let ~$\widetilde{\mathcal{H}}(v)$ denote a complete weighted graph on the vertex set $\{g_{1},\cdots,g_{m}\}$ such that the weight on any edge $\{g_{i},g_{j}\}$ is $\widetilde{c}_{v}\left(i,j\right):=\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v).$
@@
The relationship between the graphs $\mathcal{H}(v)$ and ~$\widetilde{\mathcal{H}}(v)$ stems from Lemma 13 which relates the degree of node $v$ in $G_{i}^{\prime}\wedge G_{j}^{\prime}$ to the inclusion of $v$ in $\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$ for each $i$ and $j$. Since the graph $G_{i}^{\prime}\wedge G_{j}^{\prime}\sim\mathrm{RG}(n,\mathbf{P}s^{2})$Lemma 13 implies that with probability $1-o(1/n)$, if a pair $\{g_{i},g_{j}\}$ has edge weight $\widetilde{c}_{ij}\geq k+2/\alpha$ in ~$\widetilde{\mathcal{H}}(v)$, then the corresponding edge $\{g_{i},g_{j}\}$ is present in the transitivity graph $\mathcal{H}(v)$, i.e. $v$ is correctly matched between $G_{i}$ and $G_{j}$ in the instance of pairwise $k$-core matching between them. The graph $\mathcal{H}(v)$ is not connected only if it contains a (non-empty) vertex cut $U\subset\{1,\cdots,m\}$ with no edge crossing between $U$ and $U^{c}$. For a cut $U$, let $c_{v}(U)$ denote the number of such crossing edges in $\mathcal{H}(v)$. Furthermore, define the cost of the cut $U$ in ~$\widetilde{\mathcal{H}}(v)$ as
@@
Lemma 13 is a statement about a single graph, but it can be invoked to prove the following.
Lemma 24
Let $G_{1},\cdots,G_{m}$ be correlated inhomogeneous ranodm graphs from the subsampling model with parameters $\mathbf{P}$ and $s$. Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$. Let $v\in[n]$ and let $U$ be a vertex cut of $\{1,\cdots,m\}$ such that $|U|\leq\lfloor m/2\rfloor$. Then,
($\mathbb{P}\left(\left\{c_{v}(U)=0\right\}\cap\left\{\widetilde{c}_{v}(U)>\frac{m^{2}}{4}\left(k+\frac{2}{\alpha}\right)\right\}\right)=o(1/n).$ (23)
The proof of Lemma 24 is deferred to Section 5.3. It suffices to analyze the probability that the graph ~$\widetilde{\mathcal{H}}(v)$ has a cut $U$ such that its cost $\widetilde{c}_{v}(U)$ is too small. To that end, we show that the bottleneck arises from vertex cuts of small size. Formally,
Lemma 25
Let $G_{1},\cdots,G_{m}$ be correlated inhomogeneous graphs obtained from the subsampling model. Let $v\in[n]$ and let $U_{\ell}$ denote the set $\{1,\cdots,\ell\}$ for $\ell$ in $\{1,\cdots,\lfloor m/2\rfloor\}$. For any vertex cut $U$ of $\{1,\cdots,m\}$, let $\widetilde{c}_{v}(U)$ denote its cost in the graph ~$\widetilde{\mathcal{H}}(v)$. The following stochastic ordering holds:
@@
The proof of Lemma 25 is deferred to Appendix C due to space constraints. Lemmas 24 and 25 imply that the tightest bottleneck to the connectivity of $\mathcal{H}(v)$ is the event that $\widetilde{c}_{v}(U_{1})$ is below the threshold $r:=\frac{m^{2}}{4}\left(k+\frac{2}{\alpha}\right)$, i.e. the sum of degrees of $v$ over the intersection graphs $(G_{1}\wedge G_{j}^{\prime}:j=2,\cdots,m)$ is less than $r$. This event occurs only if the degree of $v$ is simultaneously less than $r$ in each of the intersection graphs $(G_{1}\wedge G_{j}^{\prime}:j=2,\cdots,m)$. The last ingredient to the analysis is that under the condition $\sum_{i=1}^{n}e^{-d_{i}s\left(1-(1-s)^{m-1}\right)}=o(n^{-\alpha})$, this event occurs with probability $o(1/n)$.
Lemma 26
@@
Let $G_{1},\cdots,G_{m}$ be obtained from the subsampling model with parameters $\mathbf{P}$ and $s$. Let $d_{i}=\sum_{j\neq i}p_{ij}$ and $v^{*}\in\operatorname*{arg\,min}_{v\in V}d_{v}$. Let $k\geq\frac{12}{1-\alpha+2\varepsilon}$ be constant and let $r\triangleq\frac{m^{2}}{4}\left(k+\frac{2}{\alpha}\right)$. Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$ and
$\sum_{v\in V\setminus v^{*}}e^{-d_{v}s\left(1-(1-s)^{m-1}\right)}=o(n^{-\alpha}).$ (24)
Then
@@
$(\star)\triangleq\mathbb{P}\left(\exists\ v\in V\setminus v^{*}:\big{\{}\delta_{G_{1}\wedge G_{2}^{\prime}}(v)\leq r\big{\}}\cap\cdots\cap\big{\{}\delta_{G_{1}\wedge G_{m}^{\prime}}(v)\leq r\big{\}}\right)=o(1).$
Consequently, under the sufficient condition (1), the transitivity graph $\mathcal{H}(v)$ is connected for each node $v$. The above ideas are used to prove Theorem 1.
5.2 Piecing it together: Proof of Theorem 1
@@
Proof of Theorem 1. Let $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ denote the output of Algorithm 1 with $k\geq\frac{12}{1-\alpha+2\varepsilon}$. Let $E_{1}$ denote the event that Algorithm 1 fails to match all $m$ graphs exactly, i.e.
$E_{1}=\left\{\widehat{\pi}_{12}\ \neq\pi^{*}_{12}\right\}\cup\cdots\cup\left\{\widehat{\pi}_{1m}\neq\pi^{*}_{1m}\right\}.$
@@
First, we show that the output of Algorithm 1 is correct with probability $1-o(1)$ whenever the sufficient condition holds. If the event $E_{1}$ occurs, then either step 1 failed, i.e. there is a $k$-core matching $\widehat{\mu}_{ij}$ that is incorrect, or step 2 failed, i.e. at least one of the graphs $\mathcal{H}(v)$ is not connected. Therefore,
@@
$\mathbb{P}\left(E_{1}\right)$ $\leq\mathbb{P}\left(\bigcup_{i,j}\ \bigcup_{v\in\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})}\!\!\!\!\left\{\widehat{\mu}_{ij}\neq\pi^{*}_{ij}\right\}\!\right)\!+\mathbb{P}\left(\bigcup_{v\in[n]}\left\{\mathcal{H}(v)\text{ is not connected}\right\}\!\right)\!\leq o(1)+q,$
@@
where the last step uses Corollary 21, and $q$ denotes the probability that a transitivity graph $\mathcal{H}(v)$ is not connected for some $v$ in $[n]$. For each $\ell$ in the set $\{1,\cdots,\lfloor m/2\rfloor\}$, let $U_{\ell}$ denote the set $\left\{1,\cdots,\ell\right\}$. Then,
@@
$q\!$ $=\mathbb{P}\left(\bigcup_{v\in[n]}\bigcup_{\ell=1}^{\lfloor m/2\rfloor}\left\{\exists\ U\subset\{1,\cdots,m\}:|U|=\ell\text{ and }c_{v}(U)=0\right\}\right)$
@@
$\leq\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\cdot\mathbb{P}\left(\exists\ v\in[n]:c_{v}(U_{\ell})=0\right)$
@@
$\leq\!\sum_{\ell=1}^{\lfloor m/2\rfloor}m^{\ell}\Bigg{[}\mathbb{P}\left(\exists\ v\in[n]:\widetilde{c}_{v}(U_{\ell})\leq\frac{m^{2}}{4}\big{(}k+2/\alpha\big{)}\right)$
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\mathbb{P}\left(\exists\ v\in[n]:\left\{c_{v}(U_{\ell})=0\right\}\cap\big{\{}\widetilde{c}_{v}(U_{\ell})>\frac{m^{2}}{4}\big{(}k\!+\!2/\alpha\big{)}\big{\}}\right)\Bigg{]}$
@@
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!m^{\ell}\left[\mathbb{P}\left(\exists\ v\in[n]:\widetilde{c}_{v}(U_{\ell})\leq\frac{m^{2}}{4}\big{(}k\!+\!2/\alpha\big{)}\!\right)\!+\!o\left(1\right)\right]$